# The influence of cell phenotype on collective cell invasion into the extracellular matrix

### v1

@@ 1 Introduction
Collective cell invasion is a fundamental, tightly coordinated process involving both migration and proliferation, and it takes place throughout development, wound healing, and cancer metastasis [1, 2, 3, 4]. This coordinated invasion is driven by cell–cell interactions and extrinsic cues provided by the surrounding microenvironment, with the extracellular matrix (ECM) playing a pivotal role [5]. The ECM is a complex and dynamic macromolecular network composed of various proteins, such as collagen, fibronectin, and glycosaminoglycans [6]. Collagen fibres are the predominant component of the ECM, and are essential for maintaining the mechanical integrity of tissues under stress [7]. Cells actively remodel the ECM by secreting, degrading, and reorganising its components as they migrate [8]. This dynamic interplay between cells and the ECM underlies many physiological and pathological processes, however its huge complexity means that it remains challenging to fully characterise using experiments alone. An additional complicating factor is that cells display a range of phenotypes, displaying differences in, for example, in their response to their local biomechanical environment, metabolic activity, protrusive activity and connectivity patterns [9, 10, 11, 12, 13, 14].
Mathematical modelling offers a powerful approach to systematically investigate the roles of various cell–cell and cell–ECM interactions in driving collective cell migration, and to explore how differences in cell phenotype drive patterns of collective cell invasion. The overarching aim of this work is to develop a computationally tractable model of collective cell migration that includes both cell–cell interactions, such as adhesion and volume filling, and cell–ECM interactions, which lead to contact guidance and ECM re-modelling, and use it to explore how they together drive collective cell migration. In particular we focus on elucidating how differences in the cell phenotype, in terms of their ability to migrate, secrete, degrade, and interact with the ECM, impact collective cell invasion signatures.
@@ 1.1 Cell–ECM interactions during collective migration
The intricate bi-directional feedback between cells and the ECM during collective cell invasion presents a significant challenge in disentangling how specific mechanisms drive certain dynamics and how these dynamics, in turn, influence each other. In this study, we aim to gain insights through the development of a tracable mathematical model to address the overarching question of how the bi-directional interactions between cells and the ECM drive the dynamics of collective cell migration while simultaneously shaping the architecture of the ECM, and how these patterns of collective invasion are altered in cell populations with different phenotypes.
@@ 2 Mathematical model
We formulate a two-dimensional hybrid mathematical model of collective cell motility mediated by collagen fibres. Cells are modelled as discrete agents migrating over collagen fibres, and the collagen fibres are modelled via a continuous fibre field. The model incorporates the following key mechanisms: random cell motility and cell–cell interactions, both modulated by cell–fibre interactions; cell proliferation; and collagen fibre secretion and degradation by cells. We introduce the representation of cells and collagen fibres in Section 2.1, and the equations determining their evolution are given in Section 2.2. Throughout we adopt a two-dimensional Cartesian coordinate system denoted by $\mathbf{x}:=\left(x,y\right)^{\mathrm{T}}$, and $t$ denotes time. Key variables and parameters are summarised in Tables 1 and 2, respectively, and a schematic illustration of the model can be found in Figure 1.
@@
$\mathbf{u}_{\rm ave}^{i}(\mathbf{X}^{i},t;m)$ average velocity of cell $i$ during the time interval $[t-m,t]$
@@
$G$ domain of interest $[0,360]\times[0,360]$ $\mu\text{m}^{2}$
$[0,540]\times[0,540]$ $\mu\text{m}^{2}$
@@
$\Delta t$ numerical time step $1$ min
@@ 2.1 Representations for cells and collagen fibres
Here $\hat{\mathbf{u}}(\theta)=\left(\cos\theta,\sin\theta\right)^{\mathrm{T}}$ denotes a unit vector at angle $\theta$ with respect to the positive $x$-axis, and $\rho(\theta,\mathbf{x},t)\in[0,1]$ is the area fraction occupied by collagen fibres, which we use to represent collagen fibre density, with orientation $\theta$ at position $\mathbf{x}$ and time $t$. Note that collagen fibres are bidirectional [30, 31] so that $\theta\in[0,\pi)$. To enable computational tractability, we follow [33] and represent $\mathbf{\Omega}(\mathbf{x},t)$ in a diagonalised form:
@@
where $\hat{\mathbf{v}}_{1,2}(\mathbf{x},t)$ are the orthonormal eigenvectors of $\mathbf{\Omega}$ which capture the major and minor collagen fibre orientations, respectively, and $\lambda_{1}(\mathbf{x},t)\geq\lambda_{2}(\mathbf{x},t)\in[0,1]$ are the associated eigenvalues which represent the area fractions of collagen fibres in directions $\hat{\mathbf{v}}_{1}$ and $\hat{\mathbf{v}}_{2}$, respectively. Note that
$0\leq\lambda_{1}\left(\mathbf{x},t\right)+\lambda_{2}\left(\mathbf{x},t\right)=\frac{1}{\pi}\int_{0}^{\pi}\rho\left(\theta,\mathbf{x},t\right)\text{d}\theta\leq 1$ (3)
@@ 2.2.1 Cell dynamics
The choice of $\Lambda$ ensures key properties are satisfied: when the space is saturated with collagen fibres ($\lambda_{1}+\lambda_{2}=1$), $\Lambda=1$ and $\hat{\mathbf{M}}=\hat{\mathbf{\Omega}}$, and the impact of contact guidance is maximal; whereas in the absence of collagen fibres, $\Lambda=0$ and $\hat{\mathbf{M}}=\mathbf{I}$ so that the cells experience no contact guidance. Figure 1(c) visualises how $\hat{\mathbf{M}}$ reorients any $2\times 1$ vector $\mathbf{b}$, and Figure 1(d)–(e) compares the angle between $\mathbf{b}$ and $\hat{\mathbf{v}}_{1}$ (the major orientation of collagen fibres) both with and without contact guidance. As shown in Figure 1(d), $\hat{\mathbf{M}}$ automatically encodes the anisotropy information, $a$. As $a$ approaches zero (isotropic collagen fibres), $\hat{\mathbf{M}}$ tends towards the identity matrix, $\mathbf{I}$, and there is no contact guidance. Moreover, Figure 1(e) indicates that there is a switch between strong and weak contact guidance around $\lambda_{1}+\lambda_{2}=\bar{\lambda}$, as expected. Note that Figure 1(d)–(e) are symmetric around $90^{\circ}$ as collagen fibres are bi-directional.
Figure 1: Schematic illustration of the model. (a) Cells $i$ and $j$ are depicted as blue circular discs with radius $\sigma/2$ centred at $\mathbf{X}^{i}$ and $\mathbf{X}^{j}$, respectively. They sit on top of collagen fibres, where major fibre orientation $\hat{\mathbf{v}}_{1}$ and the associated area fraction $\lambda_{1}$ are visualised using brown arrows. (b) The magnitude of the pairwise interaction force is given by the function $F(|\mathbf{X}^{i}-\mathbf{X}^{j}|)$ defined in Equation (7). Orange region: short-range repulsion due to the effects of volume filling; blue region: mid-range adhesion; grey region: no cell–cell interactions. The dashed grey line represents the force arising from the Lennard-Jones potential, and $F_{0}>0$ sets the maximum repulsive force used in the model. (c) Schematic diagram of how the contact guidance matrix $\hat{\mathbf{M}}$ reorients vector $\mathbf{b}$ given vertical $\hat{\mathbf{v}}_{1}$ and horizontal $\hat{\mathbf{v}}_{2}$. (d),(e) Illustration of the mechanisms of contact guidance indicating how $\hat{\mathbf{M}}$, defined in Equation (8), reorients a vector $\mathbf{b}$, given different anisotropy degrees, $a$ (d), and total area fractions, $\lambda_{1}+\lambda_{2}$ (e). In (d), $\lambda_{1}+\lambda_{2}=1.0$ and in (e), $a=0.8$.
@@ 2.2.2 Cell proliferation
Following previous work [50], we implement a model of cell density-dependent cell proliferation in which the probability that a cell divides in the time step $[t,t+\Delta{t})$ is given by
@@ 2.2.3 Collagen fibre dynamics
We assume that cells evolve the collagen fibres through degradation and secretion. The equation of motion for collagen fibres is given by Equation (4b), where $\mathbf{g}$ is as follows:
@@
where $s>0$ and $d>0$ are the collagen secretion and degradation rates per cell, respectively. The term $\left(1-\lambda_{1}-\lambda_{2}\right)$ ensures no fibre secretion when the space is fully occupied by collagen fibres. The secretion of collagen fibres by cell $i$ is in the direction of cell $i$’s average velocity, $\mathbf{u}_{\rm ave}^{i}$, over $[t-m,t]$, where [^1] We set $\mathbf{X}^{i}(t-m)=\mathbf{X}^{i}(0)$ if $t<m$, where $\mathbf{X}^{i}(0)$ is the initial position for cell $i=1$.
@@
$\omega\left(\mathbf{X}^{i},\mathbf{x},\sigma,\omega_{0}\right)=\begin{cases}1-\frac{|\mathbf{X}^{i}-\mathbf{x}|}{\sigma/2},&\quad|\mathbf{X}^{i}-\mathbf{x}|\leq\sigma/2,\\ 0,&\quad|\mathbf{X}^{i}-\mathbf{x}|>\sigma/2.\end{cases}$ (13)
@@ 3 Computational implementation of the model
In this paper, we use two distinct two-dimensional rectangular domains, $G$, and final time, $T$, to illustrate the results presented in later Section 4:
  1. Setup 1. $G=[0,540]\times[0,540]\,\mu\mathrm{m}^{2}$ and $T=96$ hours, used to showcase the main features of the model.
  2. Setup 2. $G=[0,360]\times[0,360]\,\mu\mathrm{m}^{2}$ and $T=60$ hours, used to investigate the effects of different collagen fibre initial conditions and cell properties on dynamics.
We impose periodic boundary conditions for the cells, and the domain $G$ is discretised as $G_{\delta}$ with a grid size of $\delta=2\,\mu\mathrm{m}$, and the numerical time step is set to $\Delta{t}=1$ minute. The initial conditions are listed in Supplementary Information Section S5. The collagen fibre field, $\mathbf{\Omega}$, is discretised on $G_{\delta}$ as $\mathbf{\Omega}_{\delta}$. At each time step we
@@
8:              Let cell $i$ proliferate with probability $P_{\rm p}$ given by Equation (10)).
@@ 4 Results
We now demonstrate how differences in cell phenotypes impact patterns of collective cell invasion, underpinned by the interactions between cells and collagen fibres. We first, in Section 4.1, explore the collagen fibre distributions that arise de novo from a migratory cell phenotype, before moving to understand how different patterns of collagen fibres drive collective cell invasion in a phenotype that does not re-model the underlying fibre bed in Section 4.2. We then, in Section 4.3, simulate the dynamics of degradatory and secretory phenotypes in turn, to tease apart the subtleties of the two-way coupling between cells and collagen fibres. In Section 4.4, we conduct a systematic parameter sweep to reveal how a cell phenotype able to fully re-model collagen via both secretion of new fibres and degradation of old fibres can undergo collective invasion. Finally, in Section 4.5 we showcase the wide range of possible invasion behaviours as the relative contributions of different processes are varied.
@@ 4.1 Secreted collagen fibre dynamics reflect cell migration patterns
In Figure 2, we showcase how collective cell invasion leads to patterns in collagen fibre dynamics. We take as initial conditions $100$ cells clustered in the centre of the domain, with no collagen fibres present. As the cells migrate outwards and proliferate, they lay down collagen fibres that are oriented in the direction of outward expansion. Figure 2(a) shows snapshots of the cell positions, the total area fraction of collagen fibres ($\lambda_{1}+\lambda_{2}$), the major orientation ($\hat{\mathbf{v}}_{1}$), and the anisotropy degree ($a$), shown in grey, green, and blue shading, respectively. Cells at the leading edge exhibit directed motion due to population pressure, resulting in the secretion of highly aligned collagen fibres (Figure 2(b)). Cells behind the front are limited in their motility due to the lack of free space leading to a more isotropic collagen fibre distribution (Figure 2(c)).
Figure 2: Secreted collagen fibre dynamics reflect cell migration patterns. (a) Cell and collagen fibre distributions at $6$, $42$, and $72$ hours. Cells are visualised as blue circles with radius $\sigma/2$. The pale yellow backgrounds denote regions devoid of collagen fibres. (b) Distribution of the angle between $\hat{\mathbf{v}}_{1}$ and the positive $x$-axis in the left and right regions (LR) compared with the top and bottom regions (TB) (the regions ‘L’, ‘R’, ‘T’, and ‘B’ are indicated in the bottom-left of (a)). (c) Average anisotropy degree within concentric rings from the domain centre. Simulation details: Setup1 in Section 3. Initially, $100$ cells are arranged in a densely packed circular disc at the centre of the domain, forming a confluent arrangement, with no collagen fibres present ($\lambda_{1}+\lambda_{2}=0$). Collagen fibre secretion and degradation rates are 0.025 $0.025\text{ min}^{-1}$ and 0.0025 $0.0025\text{ min}^{-1}$, respectively.

### v2

@@ 1 Introduction
Collective cell invasion is a fundamental, tightly coordinated process involving both migration and proliferation, and it takes place throughout development, wound healing, and cancer metastasis [1, 2, 3, 4]. This coordinated invasion is driven by cell–cell interactions and extrinsic cues provided by the surrounding microenvironment, with the extracellular matrix (ECM) playing a pivotal role [5]. The ECM is a complex and dynamic macromolecular network composed of various proteins, such as collagen, fibronectin, and glycosaminoglycans [6]. Collagen fibres are the predominant component of the ECM, and are essential for maintaining the mechanical integrity of tissues under stress [7]. Cells actively remodel the ECM by secreting, degrading, and reorganising its components as they migrate [8]. This dynamic interplay between cells and the ECM underlies many physiological and pathological processes, however its huge complexity means that it remains challenging to fully characterise using experiments alone. An additional complicating factor is that cells exhibit phenotypic heterogeneity, reflected in their response to their local biomechanical environment, metabolic activity, protrusive activity and connectivity patterns [9, 10, 11, 12, 13, 14].
Mathematical modelling offers a powerful approach to systematically investigate the roles of various cell–cell and cell–ECM interactions in driving collective cell migration, and to explore how differences in cell phenotype drive patterns of collective cell invasion. The overarching aim of this work is to develop a computationally tractable model of collective cell migration that includes both cell–cell interactions, such as adhesion and volume filling, and cell–ECM interactions, which lead to contact guidance and ECM evolution, and use it to explore how they together drive collective cell migration. In particular we focus on elucidating how differences in the cell phenotype, in terms of their ability to migrate, secrete, degrade, and interact with the ECM, impact collective cell invasion signatures.
@@ 1.1 Cell–ECM interactions during collective migration
The intricate bidirectional feedback between cells and the ECM during collective cell invasion presents a significant challenge in disentangling how specific mechanisms drive certain dynamics and how these dynamics, in turn, influence each other. In this study, we aim to gain insights through the development of a tractable mathematical model to address the overarching question of how the bidirectional interactions between cells and the ECM drive the dynamics of collective cell migration while simultaneously shaping the architecture of the ECM, and how these patterns of collective invasion are altered in cell populations with different phenotypes.
@@ 2 Mathematical model
We formulate a two-dimensional hybrid mathematical model of collective cell motility mediated by collagen fibres. Cells are modelled as discrete agents migrating on top of collagen fibres, and the collagen fibres are modelled via a continuous fibre field. The model incorporates the following key mechanisms: random cell motility and cell–cell interactions, both modulated by cell–fibre interactions; cell proliferation; and collagen fibre secretion and degradation by cells. We introduce the representation of cells and collagen fibres in Section 2.1, and the equations determining their evolution are given in Section 2.2. Throughout we adopt a two-dimensional Cartesian coordinate system denoted by $\mathbf{x}:=\left(x,y\right)^{\mathrm{T}}$ with time represented by the independent variable $t$. Key variables and parameters are summarised in Tables 1 and 2, respectively, and a schematic illustration of the model can be found in Figure 1.
@@
$\mathbf{u}_{\rm ave}^{i}(\mathbf{X}^{i},t;m)$ average velocity of cell $i$ during the time interval [tm,t][t-m,t]
@@
$G$ domain of interest $0,360]\times[0,360]$ $\mu\text{m}^{2}$
$0,540]\times[0,540]$ $\mu\text{m}^{2}$
@@
$\Delta t$ numerical time step 11 min
@@ 2.1 Representations for cells and collagen fibres
Here $\hat{\mathbf{u}}(\theta)=\left(\cos\theta,\sin\theta\right)^{\mathrm{T}}$ denotes a unit vector at angle $\theta$ with respect to the positive $x$-axis, and $\rho(\theta,\mathbf{x},t)\in[0,1]$ is the area fraction occupied by collagen fibres, which we use to represent collagen fibre density, with orientation $\theta$ at position $\mathbf{x}$ and time $t$. Note that collagen fibres are bidirectional [^1] This means that the two orientations of a fibre, separated by $180^{\circ}$, are indistinguishable. [30, 31] so that $\theta\in[0,\pi)$. To enable computational tractability, we follow [33] and represent $\mathbf{\Omega}(\mathbf{x},t)$ in a diagonalised form:
@@
where $\hat{\mathbf{v}}_{1,2}(\mathbf{x},t)$ are the orthonormal eigenvectors of $\mathbf{\Omega}$ which capture the major and minor collagen fibre orientations, respectively, and $\lambda_{1}(\mathbf{x},t)\geq\lambda_{2}(\mathbf{x},t)\in[0,1]$ are the associated eigenvalues which represent the area fractions of collagen fibres in directions $\hat{\mathbf{v}}_{1}$ and $\hat{\mathbf{v}}_{2}$, respectively. Note that for the total area fraction, which represents collagen fibre density, we have
$0\leq\lambda_{1}\left(\mathbf{x},t\right)+\lambda_{2}\left(\mathbf{x},t\right)=\frac{1}{\pi}\int_{0}^{\pi}\rho\left(\theta,\mathbf{x},t\right)\text{d}\theta\leq 1.$ (3)
@@ 2.2.1 Cell dynamics
The choice of $\Lambda$ ensures key properties are satisfied: when the space is saturated with collagen fibres ($\lambda_{1}+\lambda_{2}=1$), $\Lambda=1$ and $\hat{\mathbf{M}}=\hat{\mathbf{\Omega}}$, and the impact of contact guidance is maximal; whereas in the absence of collagen fibres, $\Lambda=0$ and $\hat{\mathbf{M}}=\mathbf{I}$ so that the cells experience no contact guidance. Figure 1(c) visualises how $\hat{\mathbf{M}}$ reorients any $2\times 1$ vector $\mathbf{b}$, and Figure 1(d)–(e) compares the angle between $\mathbf{b}$ and $\hat{\mathbf{v}}_{1}$ (the major orientation of collagen fibres) both with and without contact guidance. As shown in Figure 1(d), $\hat{\mathbf{M}}$ automatically encodes the anisotropy information $a$. As $a$ approaches zero (isotropic collagen fibres), $\hat{\mathbf{M}}$ tends towards the identity matrix, $\mathbf{I}$, and there is no contact guidance. Moreover, Figure 1(e) indicates that there is a switch between strong and weak contact guidance around $\lambda_{1}+\lambda_{2}=\bar{\lambda}$, as expected. Note that Figure 1(d)–(e) are symmetric around $90^{\circ}$ because collagen fibres are bidirectional. Figure 1 also illustrates the advantage of representing the collagen fibres as a tensorial field: the matrix-vector multiplication ($\hat{\mathbf{M}}\mathbf{b}$) directly reorients the cell’s migratory direction without the need for additional assumptions to select between the two directions (see Supplementary Information Section S1), due to the bidirectionality of collagen fibres.
Figure 1: Schematic illustration of the model. (a) Cells $i$ and $j$ are depicted as blue circular discs with radius $\sigma/2$ centred at $\mathbf{X}^{i}$ and $\mathbf{X}^{j}$, respectively. They sit on top of collagen fibres, where major fibre orientation $\hat{\mathbf{v}}_{1}$ and the associated area fraction $\lambda_{1}$ are visualised using brown arrows. (b) The magnitude of the pairwise interaction force is given by the function $F(|\mathbf{X}^{i}-\mathbf{X}^{j}|)$ defined in Equation (7). Orange region: short-range repulsion due to the effects of volume filling; blue region: mid-range adhesion; grey region: no cell–cell interactions. The dashed grey line represents the force arising from the Lennard-Jones potential, and $F_{0}>0$ sets the maximum repulsive force used in the model. (c) Schematic diagram of how the contact guidance matrix $\hat{\mathbf{M}}$ reorients vector $\mathbf{b}$ given vertical $\hat{\mathbf{v}}_{1}$ and horizontal $\hat{\mathbf{v}}_{2}$. (d),(e) Illustration of the mechanisms of contact guidance indicating how $\hat{\mathbf{M}}$, defined in Equation (8), reorients a vector $\mathbf{b}$, given different anisotropy degrees, $a$ (d), and total area fractions, $\lambda_{1}+\lambda_{2}$ (e). In (d), $\lambda_{1}+\lambda_{2}=1.0$ and in (e), $a=0.8$. In (e), the arrow indicates the strength of contact guidance at the critical total fibre area fraction $\bar{\lambda}=0.4$.
@@ 2.2.2 Cell proliferation
Following previous work [50], we implement a model of cell density-dependent cell proliferation in which the probability that a cell divides in the time step $t,t+\Delta{t})$ is given by
@@ 2.2.3 Collagen fibre dynamics
We assume that cells evolve the collagen fibres through degradation and secretion. The governing equation for collagen fibres is given by Equation (4b), where $\mathbf{g}$ is as follows:
@@
where $s>0$ and $d>0$ are the collagen secretion and degradation rates per cell, respectively. The term $\left(1-\lambda_{1}-\lambda_{2}\right)$ ensures no fibre secretion when the space is fully occupied by collagen fibres. The secretion of collagen fibres by cell $i$ is in the direction of cell $i$’s average velocity, $\mathbf{u}_{\rm ave}^{i}$, over [tm,t][t-m,t], where [^2] We set $\mathbf{X}^{i}(t-m)=\mathbf{X}^{i}(0)$ if $t<m$, where $\mathbf{X}^{i}(0)$ is the initial position for cell $i=1$.
@@
$\omega\left(\mathbf{X}^{i},\mathbf{x},\sigma\right)=\begin{cases}1-\frac{|\mathbf{X}^{i}-\mathbf{x}|}{\sigma/2},&\quad|\mathbf{X}^{i}-\mathbf{x}|\leq\sigma/2,\\ 0,&\quad|\mathbf{X}^{i}-\mathbf{x}|>\sigma/2.\end{cases}$ (13)
@@ 3 Computational implementation of the model
In this paper, we use two distinct two-dimensional rectangular domains, $G$, and final times, $T$, to illustrate the results presented in later Section 4:
  1. Setup 1. $G=[0,540]\times[0,540]\,\mu\mathrm{m}^{2}$ and $T=96$ hours, used to showcase the main features of the model in detail.
  2. Setup 2. $G=[0,360]\times[0,360]\,\mu\mathrm{m}^{2}$ and $T=60$ hours (smaller domain size and longer simulation time for computational efficiency), used to investigate the effects of different collagen fibre initial conditions and cell properties on dynamics.
We impose periodic boundary conditions for the cells, and the domain $G$ is discretised as $G_{\delta}$ with a grid size of $\delta=2\,\mu\mathrm{m}$, and the numerical time step is set to $\Delta{t}=1$ minute. The initial conditions are listed in Supplementary Information Section S5. The collagen fibre field, $\mathbf{\Omega}$, is discretised on $G_{\delta}$ as $\mathbf{\Omega}_{\delta}$. At each time step we:
@@
8:     Let cell $i$ proliferate with probability $P_{\rm p}$ given by Equation (10).
@@ 4 Results
We now demonstrate how differences in cell phenotypes impact patterns of collective cell invasion, underpinned by the interactions between cells and collagen fibres. We first, in Section 4.1, explore the collagen fibre distributions that arise de novo from a migratory cell phenotype, before moving to understand how different patterns of collagen fibres drive collective cell invasion in a phenotype that does not re-model the underlying fibre bed in Section 4.2. We then, in Section 4.3, simulate the dynamics of degrading and secreting phenotypes in turn, to tease apart the subtleties of the two-way coupling between cells and collagen fibres. In Section 4.4, we conduct a systematic parameter sweep to reveal how a cell phenotype able to fully re-model collagen via both secretion of new fibres and degradation of old fibres can undergo collective invasion. Finally, in Section 4.5 we showcase the wide range of possible invasion behaviours as the relative contributions of different processes are varied. All cell phenotypes and their corresponding dynamics explored are listed in Table 3.
Cell phenotypes Section
Non-collagen modulating phenotype Section 4.2
Collagen degrading phenotype Section 4.3
Collagen secreting phenotype
Collagen-modulating phenotype Section 4.4
Cell-cell interaction migratory phenotype Section 4.5
Random motility migratory phenotype
Contact guidance (in)sensitive phenotype(s)
Short (& long) memory phenotype(s)
Table 3: Cell phenotypes and their corresponding dynamics explored in Section 4.
@@ 4.1 Secreted collagen fibre dynamics reflect cell migration patterns
In Figure 2, we showcase how collective cell invasion leads to patterns in collagen fibre dynamics. We take as initial conditions $100$ cells clustered in the centre of the domain, with no collagen fibres present. As the cells migrate outwards and proliferate, they lay down collagen fibres that are oriented in the direction of outward expansion. Figure 2(a) shows snapshots of the cell positions, the total area fraction of collagen fibres ($\lambda_{1}+\lambda_{2}$), the major orientation ($\hat{\mathbf{v}}_{1}$), and the anisotropy degree ($a$), shown in grey, green, and blue shading, respectively. Cells at the leading edge exhibit directed motion due to population pressure, resulting in the secretion of highly aligned collagen fibres (Figure 2(b)). Cells behind the front are limited in their motility due to the lack of free space leading to a more isotropic collagen fibre distribution (Figure 2(c)). We note that the dynamics of cells and collagen fibres are qualitatively insensitive to the functional form of the collagen-modulating kernel, $\omega$ (see Figures S4S5).
Figure 2: Secreted collagen fibre dynamics reflect cell migration patterns. (a) Cell and collagen fibre distributions at $6$, $42$, and $72$ hours. Cells are visualised as blue circles with radius $\sigma/2$. The pale yellow backgrounds denote regions devoid of collagen fibres. (b) Distribution of the angle between $\hat{\mathbf{v}}_{1}$ and the positive $x$-axis in the left and right regions (LR) compared with the top and bottom regions (TB) (the regions ‘L’, ‘R’, ‘T’, and ‘B’ are indicated in the bottom-left of (a)). The violin plot shows the kernel density estimate of this distribution, with a miniature box plot overlaid: the thick black bar represents the interquartile range (IQR, Q1–Q3), the white dot indicates the median, and the thin black lines (‘whiskers’) extend to data points within $1.5\timesIQR$ from Q1 and Q3. (c) Average anisotropy degree within concentric rings from the domain centre. Simulation details: Setup1 in Section 3. Initially, $100$ cells are arranged in a densely packed circular disc at the centre of the domain, forming a confluent arrangement, with no collagen fibres present ($\lambda_{1}+\lambda_{2}=0$). Collagen fibre secretion and degradation rates are $0.025\text{ min}^{-1}$ and $0.0025\text{ min}^{-1}$, respectively.